Acct 2270 V0852 2018WI (B) Kyle Maisano 3/16/18 10:41 PM Test: Weeh #1 Chapter #2 Test Time Remaining: 00:53:40 Submit Test This Question: 1 pt Alkinson Company wishes to ean after-tax net income of $18,000. Total fixed costs are $84,000 and the contribution margin is $8.00 per unt Asinsons tox roateis 40%. The number of units that must be sold to earn the targeted net income is O A. 17,000 O B. 21,500 O c. 19,000 OD, 14,000 Acct 2270 v0852 2018WI (Bp is bssed on
Answer #1
C 19,000
Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
Paticulars Amount
After Tax net Income              18,000.00
Tax rate 40%
Before tax net income = 18000/60%              30,000.00
Fixed costs              84,000.00
Contribution = 30000 + 84000            114,000.00
Contribution margin per unit                         6.00
No of units to be sold = 114,000/6              19,000.00